OpenStudy (anonymous):

What is the value of the expression (–3)*4? A. –12 B. 81 C. 12 D. –81

thomaster (thomaster):

it means -3-3-3-3

OpenStudy (anonymous):

so it whould be

OpenStudy (anonymous):

hold up let me do the math

OpenStudy (anonymous):

is it - or +

OpenStudy (the_fizicx99):

It's negative.

OpenStudy (anonymous):

k so hold up

OpenStudy (studygurl14):

It is...(-3)(-3)(-3)(-3)

OpenStudy (studygurl14):

wait is it times or to the power to?

OpenStudy (anonymous):

A right

OpenStudy (studygurl14):

If multiply: -12 If to the power to: 81

OpenStudy (the_fizicx99):

It's basically a copy of -3 four times. You can think of it as 3 * 4 and add the negative sign ( a negative and positive is negative) Or add -3 four times. -3 - 3 -3 -3

OpenStudy (anonymous):

so im right?

OpenStudy (the_fizicx99):

Yeah, seeing that there's 81 in the options. Is it being raised to a power? \[-3^{4}\]

thomaster (thomaster):

the 81 is just a deceiving trick. The * symbol means multiply so -3 times 4 -3-3-3-3 = -12

OpenStudy (anonymous):

What is the value of the expression –5x4 when x = –2? A. –40 B. –80 C. 40 D. 80

OpenStudy (studygurl14):

Well, it depends on the actual equation. What does the book actually say @icrashjk ? \[(-3)^{4}\] OR \[(-3) * 4\]

thomaster (thomaster):

\(\sf\Large -5x^4\ ?\)

OpenStudy (anonymous):

unit assement theres no book

OpenStudy (the_fizicx99):

I'm aware of that. But if it was a picture of -3^4 and he put -3*4 then it would be wrong..

OpenStudy (studygurl14):

well, you're getting it from somewhere...

OpenStudy (anonymous):

im in k12

OpenStudy (studygurl14):

well, whatever, wherever you're getting it from, what does it say?

OpenStudy (the_fizicx99):

That's irrelevant. Were asking about the question is it. \[-3\times4 OR -3^{4}\]

OpenStudy (studygurl14):

yeah!

OpenStudy (the_fizicx99):

LaTeX tabs kinda sucks >.>

OpenStudy (anonymous):

-5x*4

OpenStudy (studygurl14):

so, to confirm, you're saying multiply, right?

OpenStudy (the_fizicx99):

._. Kay. Done~ I'm just going with -5x*4 Plug in -2 for x -5(-2) * 4 -5 * -2 = ?? ?? * 4 = ??

thomaster (thomaster):

so (-5)*(-2)*4 (-5)*(-2) = 10 since multiplying 2 negative numbers will give you a positive number then just multiply 10 times 4

OpenStudy (anonymous):

40

OpenStudy (the_fizicx99):

Yes, -2 * -5 = 10 10 * 4 = 40

OpenStudy (studygurl14):

*head in hands* See ya! Good luck @tHe_FiZiCx99

OpenStudy (anonymous):

C?

OpenStudy (the_fizicx99):

Head in hands? xD wut?